OpenStudy (anonymous):

Find the sum of the series 1/(1×2) + 1/ (2×3)+ 1/(3×4)+...+1/(n×(n+1))+...

OpenStudy (zarkon):

\[\frac{1}{n(n+1)}=\frac{1}{n}-\frac{1}{n+1}\]